# A hybrid numerical method for elastic wave propagation in discontinuous media with complex geometry

### v1

@@ 4.3 Complex geometry
Finally, to demonstrate stability we consider a layered domain $\Omega=[0,1]\times[0,2]$, and let the DG domain $\Omega_{v}=[0,1]\times[0,0.2]$ contain a complex internal geometry. Moreover, we let the overall domain be discontinuous with material parameters $\mu_{u}=0.1,\lambda_{u}=1$, $\mu_{v}=0.01$, $\lambda_{v}=0.1$ and densities $\rho_{u}=\rho_{v}=1$. We also let the two meshes along the interface be conforming. The fields are then initialized at $t=0$ as two symmetric Gaussian pulses given by
@@
We then let the pules propagate freely without any external load, imposing traction free conditions on the domain boundaries as well as the internal cavities. The resulting distribution of the magnitude of the pulse at various time steps are shown in Figure 6(a)-6(d).
In Figure 6(a), we see the internal geometry in $\Omega_{v}$ clearly, which would be impossible to resolve using a curvilinear grid such as in the SBP-FD case. Moreover, we see the effect of the pulse propagating from $\Omega_{v}$ into $\Omega_{u}$ continuously. In Figure 6(b), we can clearly see the shift in wavelength of the solution as it enters a domain with larger wave speed, together with the reflections arising at the interface. Figure 6(c) then demonstrates how the wave has interacted with the northern boundary and produce reflections close to the interface. Finally at $T=100$ in Figure 6(d), we observe that the discretization remains stable.
@@ Appendix A Proof of positive definiteness
$\mathcal{Z}_{12}^{T}\mathcal{Z}_{22}^{-1}\mathcal{Z}_{12}=\begin{bmatrix}\frac{\lambda_{u}^{2}}{2\mu_{u}+\lambda_{u}}(C^{\prime}-\frac{1}{4\beta_{u}})(H_{x}\otimes I_{y})&&0\\ 0&&\mu[C^{\prime}-\frac{\mu-U}{4\beta_{u}(2\mu_{u}+\lambda_{u})}](H_{x}\otimes I_{y})\end{bmatrix}.$
Upon subtracting the above term from $\mathcal{Z}_{11}$ we note that the only remaining non-zero term of the complement is $[\mathcal{Z}/\mathcal{Z}_{22}]_{11}$. Hence, to show that $\mathcal{Z}/\mathcal{Z}_{22}$ is positive semi-definite, we need to make sure that the eigenvalues of the remaining term is non-negative. It should be noted that the term only consists of a constant which scales $H_{x}\otimes I_{y}$, thus it is sufficient to show that this constant is non-negative. Multiplying by $2\mu_{u}+\lambda_{u}$, the constant term satisfies
@@
Note that the inequality follows from the assumption of $\mu_{u}+\lambda_{u}>0$ and $\mu_{u}>0$, hence the Schur conditions are satisfied.

### v2

@@ 4.3 Complex geometry
Finally, to demonstrate stability we consider a layered domain $\Omega=[0,1]\times[0,2]$ containing a complex geometric feature. The two layers are given by $\Omega_{u}=[0,1]\times[0.2,2]$, $\Omega_{v}=[0,1]\times[0,0.2]$, where the geometric feature is in $\Omega_{v}$. Moreover, we let the material parameters in the two domains be given by $\mu_{u}=0.1,\lambda_{u}=1$, $\mu_{v}=0.01$, $\lambda_{v}=0.1$ and densities $\rho_{u}=\rho_{v}=1$. This motivates using an unstructured method in $\Omega_{v}$. We let the two meshes along the interface be conforming and use the same penalization parameters $\alpha=\beta=60$ in both domains. The fields are then initialized at $t=0$ as two symmetric Gaussian pulses given by
@@
The pules then propagate freely without any external load, imposing traction free conditions on the domain boundaries as well as the internal cavities. The resulting distribution of the magnitude of the pulse at various time steps are shown in Figure 7(a)-7(d).
In Figure 7(a), we see the internal geometry in $\Omega_{v}$ clearly. Resolving the geometry using a curvilinear FD grid would be impractical as it would require us to partition the domain into a large number of blocks to resolve the arcs and corners properly. Moreover, we see the effect of the pulse propagating from $\Omega_{v}$ into $\Omega_{u}$ continuously. In Figure 7(b), we can clearly see the shift in wavelength of the solution as it enters a domain with larger wave speed, together with the reflections arising at the interface. Figure 7(c) then demonstrates how the wave has interacted with the top boundary and produce reflections close to the interface. Finally at $T=100$ in Figure 7(d), we observe that the discretization remains stable.
@@ 5 Conclusion
Acknowledgement
A.G is thankful for the economical support in the form of a travel grant from the SJCKMS foundation.
@@ Appendix A Proof of positive definiteness
$\mathcal{Z}_{12}^{T}\mathcal{Z}_{22}^{-1}\mathcal{Z}_{12}=\begin{bmatrix}\frac{\lambda_{u}^{2}}{2\mu_{u}+\lambda_{u}}(C^{\prime}-\frac{1}{4\beta_{u}})(H_{x}\otimes I_{y})&&0\\ 0&&\mu[C^{\prime}-\frac{\mu_{u}}{4\beta_{u}(2\mu_{u}+\lambda_{u})}](H_{x}\otimes I_{y})\end{bmatrix}.$
Upon subtracting the above term from $\mathcal{Z}_{11}$ we note that the only remaining non-zero term of the complement is $[\mathcal{Z}/\mathcal{Z}_{22}]_{11}$. Hence, to show that $\mathcal{Z}/\mathcal{Z}_{22}$ is symmetric positive semi-definite, we need to make sure that the eigenvalues of the remaining term is non-negative. It should be noted that the term only consists of a constant which scales $H_{x}\otimes I_{y}$, thus it is sufficient to show that this constant is non-negative. Multiplying by $2\mu_{u}+\lambda_{u}$, the constant term satisfies
@@
Note that the inequality follows from the assumption of the Lamé parameters being positive, hence the Schur conditions are satisfied.